Gday team,
Rat testing has been put back in place for January so there is now the 0400 scheduled run again to take security over to flne, see email.
11-1 Nathan removed and replaced exhaust lagging on port main to stop the white powder being sucked up blocking the air filter. New air filter have been fitted to both mains.
Helm throttle levers are offset slightly for some reason. At idle the mains are at 650, when clutched in the port jumps to 680 but starboard stays at 650. MRF has been raised and Brodie is going to change out levers at some stage as he believes they could be the issue.
Buttons on jet panel have no identification which has been raised as a hazard with new skippers training so new buttons have been ordered which will be changed out when they arrive.
Fire equipment due for testing see flagged email
Genset strainer done 17-1
test and tag done 17-1
have a good swing